# Quantum Origin of Classical Background Fields from Coherent States

### v1

@@ 3.2 Transition to the Heisenberg Picture
We now focus on Eqs. (1921) and attempt to rewrite the formulation from the Schrödinger picture into the Heisenberg picture. As declared earlier, at the reference time $t_{0}$, the quantum states and operators in both pictures are identified [Itzykson-Zuber1980, Sakurai1967]. Thus
@@
In the Heisenberg picture, constructing the Heisenberg equation of motion is essential. This requires rewriting the Hamiltonian in Heisenberg picture. Using the relation for $\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})$, we define
@@
consisting only Heisenberg field operators. We can identify $\hat{U}(t,t_{0})$ and $\hat{U}_{\mathrm{S}}(t,t_{0})$: substitute $\hat{O}_{\mathrm{S}}=\hat{U}_{\mathrm{S}}(t,t_{0})$ into Eq. (26). This allows the time evolution operator including the $\mathcal{A}$-effects to be applied consistently in both the Heisenberg and Schrödinger pictures.
@@ 4 Furry picture
In this section, we construct the Furry picture [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009] commonly employed in strong-field QED starting from the Heisenberg picture formulation developed in the previous section. By formulating the picture dependence of the classical background field $\mathcal{A}$ in terms of coherent states and displacement operators, we have established a consistent framework in which the spacetime dependence of $\mathcal{A}$ used in standard strong-field QED calculations is well defined and justified.
@@
In this work, laser depletion is instead addressed using scattering calculations formulated within the Schrödinger–Heisenberg framework developed in Sect. 3, with explicit applications presented in Sects. 5 and 7. For comparison, related approaches based on alternative picture choices, such as those developed by Ref. [Ilderton-Seipt2018], will also be discussed. These frameworks allow the dynamical modification of the background field to be treated consistently and go beyond the standard Furry picture description. While we illustrate the construction using the Dirac field for concreteness, the same transformation applies equally to the EM field. Section 4.2 then shows how this general framework reproduces the familiar Volkov solutions of the Dirac field when the background field is taken to be a plane wave.
@@ 4.1 Basic structures
See also Sect. 4.2 about Eq. (38). We do the same for $\hat{A}_{\mathrm{H}}^{\mu}$, the field equation of $\hat{A}_{\mathrm{F}}^{\mu}$ is imposed.
@@
This formalism defines the Furry picture [Ritus1985, Piazza2012, Fedotov2023, Furry1951, Moortgat-Pick2009]. Use of $\hat{S}_{\mathrm{F}}$ allows the computation of scattering probabilities, for instance, nonlinear Compton scattering [Ritus1985, Brown-Kibble1964, Nikishov-Ritus1964, Nikishov-Ritus1967] and nonlinear Breit–Wheeler processes [Ritus1985, Nikishov-Ritus1964, Nikishov-Ritus1967, Reiss1962], as confirmed in numerous previous studies.
@@ 4.2 Volkov Solutions
We now focus on the Dirac field equation (Eq. (38)) in the Furry picture. Taking e=|e|-e=-|e| and applying the explicit form of the Hamiltonian $\hat{H}_{\mathrm{F}}^{0}(t)$, Eq. (38) becomes
$i\hbar\gamma^{\mu}\bigg[\partial_{\mu}-i\frac{e}{\hbar}\mathcal{A}_{\mu}(x)\bigg]\hat{\psi}_{\mathrm{F}}(x)-mc\hat{\psi}_{\mathrm{F}}(x)$ $=0$ (44)
When the coherent laser EM field is a plane-wave pulse, the Dirac field $\hat{\psi}_{\mathrm{F}}(x)$ admits an analytic solution in terms of the Volkov solutions [Joachain_Kylstra_Potvliege_2011, Ritus1985, Piazza2012, Fedotov2023, Volkov1936, Itzykson-Zuber1980]. In this representation, the effect of the laser background is absorbed into the Dirac field itself, enabling scattering calculations to be performed with the laser influence incorporated from the outset.
@@ 5.1 Heisenberg-Euler vacuum model
$Z[\mathcal{A}]=e^{iW[\mathcal{A}]/\hbar}$ $=\langle 0_{\mathrm{F}}|\hat{V}(\infty,\infty)|0_{\mathrm{F}}\rangle,$ (45)
where, $W[\mathcal{A}]$ is the effective action whose integrand is the HE Lagrangian density [Greiner_Muller_Rafelski_1985, Fedotov2023, Ritus1985, Schwinger1951, Heisenberg-Euler1936, Weisskopf1936, Dunne2005]. Within the present approach, this expression can be reformulated in terms of a coherent state. We reconsider Eq. (8) in the Schrödinger picture. Suppose that the system contains only a coherent state $|\alpha_{\mathrm{S}}(t)\rangle$ at all times. We then examine the corresponding $\alpha$-$\alpha$ transition amplitude:
@@
Here, $C[\mathcal{A}]$ denotes the overall c-number phase generated by the action of the background-dressed free Hamiltonian on the vacuum $|0_{\mathrm{F}}\rangle$, i.e., the part of the Hamiltonian that governs vacuum propagation in the presence of the classical background but does not generate particle excitations, analogous to the vacuum energy phase in the Gell-Mann–Low theorem. This phase does not contribute to the Heisenberg–Euler effective action and drops out of all physical observables. A conceptual advantage of the coherent state formulation is that a background field automatically satisfies the Lorenz gauge condition $\partial_{\mu}\mathcal{A}^{\mu}(x)=0$ through the Gupta–Bleuler condition applied to the underlying photon coherent state.
The goal of this subsection is not to alter the value of the HE Lagrangian, but rather to clarify its physical origin within strong-field QED when classical background fields are understood as limits of quantum-coherent EM states.
@@ 5.2 Beyond Furry picture
The present method also allows us to describe depletion effects of a coherent laser EM field. In the Furry picture, quantum transitions are treated under a fixed coherent state $\alpha$. Depletion effects corresponds to quantum processes in which a coherent state $\alpha$ evolves into another coherent state $\alpha^{\prime}$. In the Schrödinger picture, the corresponding transition amplitude is
@@
An equivalent form can be found in the work of Ref. [Fedotov2023, Ilderton-Seipt2018]. The automatic inclusion of the Lorenz gauge condition is the advantage of the present method. In the present formulation, backreaction is not introduced as an independent mechanism, but appears automatically as the response of matter fields to transitions between coherent states (depletion).
@@ 6 Connection to BRST quantization
This section clarifies the relation between our coherent-state formulation and the BRST quantization of gauge theories [BRS1976, Tyutin1975], ensuring the consistency of the physical-state conditions. The present formulation has been developed using the Gupta–Bleuler condition [Gupta1950, Bleuler1950], which provides a convenient and transparent implementation of gauge constraints in covariant QED. It is worth emphasizing, however, that this choice does not restrict the generality of the construction. In fact, the Gupta–Bleuler condition may be viewed as a special realization of the BRST formalism in the context of abelian gauge theory, where ghost degrees of freedom decouple from the physical sector (while the operator $\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}$ itself is not a Noether charge, the corresponding BRST charge provides the conserved generator that systematizes this condition into a cohomological definition of physical states).
@@
with equivalence classes taken modulo BRST-exact states. In QED, owing to the decoupling of ghost degrees of freedom, this condition reduces effectively to the Gupta–Bleuler constraint on the positive-frequency part of the gauge field [Kugo-Ojima1979, Nakanishi-Ojima1990]. The coherent states considered in this work may therefore be equivalently characterized as BRST-physical states, as long as the displacement operator preserves the BRST cohomology.
@@
7 Path integral representation
In this section, we reformulate the coherent-field QED amplitudes in the path integral representation, showing how the standard background-field generating functional emerges from coherent-state boundary conditions. Throughout this section, $\mathcal{A}$ is treated as a classical field whose status depends on boundary conditions: it is fixed in the Furry picture limit, but becomes a dynamical integration variable when depletion is included. We rewrite the scattering amplitude
@@ 7 Path integral representation
introduced in Eq. (47) as an overlap of asymptotic in–out states in the Heisenberg picture and translate it into a path integral representation. Here, $\alpha_{(\mathrm{in})}$ and $\alpha_{(\mathrm{out})}^{\prime}$ denote coherent states of the EM field associated with the laser modes at $t\to$-$\infty$ and $t\to+\infty$, respectively. The states $\Psi_{(\mathrm{in})}$ and $\Psi_{(\mathrm{out})}^{\prime}$ represent general multiparticle states, including electrons, positrons and photons other than the laser modes. All states are understood as asymptotic in/out states in the Heisenberg picture.
The starting point is the observation made in the depletion analysis [Fedotov2023, Ilderton-Seipt2018]: the amplitude $S_{fi}$ corresponds to a fully quantum transition between two distinct coherent photon states, accompanied by arbitrary changes in the remaining quantum degrees of freedom. This amplitude is therefore more general than the generating functional in a fixed background field, $Z[\mathcal{A}]$, employed in conventional strong-field QED, corresponding to the fixed-background limit of coherent-field QED which implicitly assumes a fixed background field and hence $\alpha^{\prime}=\alpha$.
To obtain a path integral representation, recall Eq. (7), we separate the total gauge field into two components,
@@
$\mathcal{A}_{(-\infty)}^{\mu}(x)=\lim_{t\to-\infty}\mathcal{A}^{\mu}(x)$ $\sim\langle\alpha_{(\mathrm{in})\mathrm{H}}|\hat{A}_{(\mathrm{in})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{in})\mathrm{H}}\rangle,$ (51)
$\mathcal{A}_{(+\infty)}^{\mu}(x)=\lim_{t\to+\infty}\mathcal{A}^{\mu}(x)$ $\sim\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}|\hat{A}_{(\mathrm{out})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}\rangle,$ (52)
for the asymptotic quantized EM filed $\hat{A}_{(\mathrm{in/out})\mathrm{H}}^{\mu}(x)$. In contrast, the quantied EM field other than the laser field and the Dirac field are treated with standard in–out boundary conditions.
@@
$\hskip 0.0pt\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ $=\int_{\mathcal{A}_{(-\infty)}}^{\mathcal{A}_{(+\infty)}}\mathcal{D}\mathcal{A}\int\mathcal{D}A\int\mathcal{D}\psi\mathcal{D}\bar{\psi}$
@@
where the functional integration over $\mathcal{A}$ is restricted to field configurations consistent with the coherent state boundary conditions specified above. The Lagrangian density $\mathcal{L}_{\mathrm{QED}}$ is related in the standard way to the Hamiltonian density $\mathcal{H}_{\mathrm{QED}}$ employed in the previous operator-based formulation. This expression provides a first-principles path integral formulation of laser depletion and backreaction, without double counting of photon degrees of freedom.
The conventional generating functional of strong-field QED [Fedotov2023] is recovered as a special case. If the laser field is treated as fixed and undepleted, so that $\alpha^{\prime}=\alpha$ and the functional integration over $\mathcal{A}$ is suppressed, the above expression reduces to
$\hskip 0.0ptZ[\mathcal{A}]$ $=\int\mathcal{D}A\int\mathcal{D}\psi\mathcal{D}\bar{\psi}\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (54)
which is precisely the generating functional used in the Furry picture formulation and in the derivation of the Heisenberg–Euler effective action [Greiner_Muller_Rafelski_1985, Schwinger1951]. From this viewpoint, $Z[\mathcal{A}]$ corresponds to a semiclassical limit of the more general coherent state transition amplitude, in which the laser field is frozen to its expectation value and quantum depletion effects are neglected.
We emphasize that the path integral representation introduced here should be understood as a rewriting of operator-level matrix elements between physical (Gupta–Bleuler [Gupta1950, Bleuler1950]/BRST [BRS1976, Tyutin1975, Kugo-Ojima1979, Nakanishi-Ojima1990]) states. Gauge constraints are therefore enforced prior to taking the functional integral, and the resulting expression does not represent an independent quantization procedure.

### v2

@@ 3.2 Transition to the Heisenberg Picture
We now focus on Eqs. (1921) and attempt to rewrite the formulation from the Schrödinger picture into the Heisenberg picture. As declared earlier, at the reference time $t_{0}$, the quantum states and operators in both pictures are identified [20, 33]. Thus
@@
In the Heisenberg picture, constructing the Heisenberg equation of motion is essential. This requires rewriting the Hamiltonian in the Heisenberg picture. Using the relation for $\hat{O}_{\mathrm{H}}(t,\boldsymbol{x})$, we define
@@
consisting only of Heisenberg field operators. We can identify $\hat{U}(t,t_{0})$ and $\hat{U}_{\mathrm{S}}(t,t_{0})$: substitute $\hat{O}_{\mathrm{S}}=\hat{U}_{\mathrm{S}}(t,t_{0})$ into Eq. (27). This allows the time evolution operator including the $\mathcal{A}$-effects to be applied consistently in both the Heisenberg and Schrödinger pictures.
@@ 4 Furry picture
In this section, we construct the Furry picture [31, 4, 6, 9, 25] commonly employed in strong-field QED (or nonlinear QED) starting from the Heisenberg picture formulation developed in the previous section. By formulating the picture dependence of the classical background field $\mathcal{A}$ in terms of coherent states and displacement operators, we have established a consistent framework in which the spacetime dependence of $\mathcal{A}$ used in standard strong-field QED calculations is well defined and justified. In particular, while textbook formulation of QED with prescribed external fields postulate a time-dependent c-number background from the outset, here the same time dependence emerges from picture choice rather than being introduced ad hoc at the Hamiltonian level, providing a conceptual link to the standard unstable-vacuum formalism [7].
@@
In this work, laser depletion is instead addressed using scattering calculations formulated within the Schrödinger–Heisenberg framework developed in Sect. 3, with explicit applications presented in Sects. 5 and 7. For comparison, related approaches based on alternative picture choices, such as those developed by Ref. [19], will also be discussed. These frameworks allow the dynamical modification of the background field to be treated consistently and go beyond the standard Furry picture description. While we illustrate the construction using the Dirac field for concreteness, the same transformation applies equally to the EM field. Section 4.2 then shows how this general framework reproduces the familiar Volkov solutions of the Dirac field when the background field is taken to be a plane wave.
@@ 4.1 Basic structures
See also Sect. 4.2 about Eq. (39). Applying the same transformation for $\hat{A}_{\mathrm{H}}^{\mu}$ yields the field equation of $\hat{A}_{\mathrm{F}}^{\mu}$.
@@
This formalism defines the Furry picture [31, 4, 6, 9, 25]. Use of $\hat{S}_{\mathrm{F}}$ allows the computation of scattering probabilities, for instance, nonlinear Compton scattering [31, 3, 27, 28] and nonlinear Breit–Wheeler processes [31, 27, 28, 30], as confirmed in numerous previous studies.
@@ 4.2 Volkov Solutions
We now focus on the Dirac field equation (Eq. (39)) in the Furry picture. Taking the electron charge to be e=|e|-e=-|e| and applying the explicit form of the Hamiltonian $\hat{H}_{\mathrm{F}}^{0}(t)$, Eq. (39) becomes
$i\hbar\gamma^{\mu}\bigg[\partial_{\mu}-i\frac{e}{\hbar}\mathcal{A}_{\mu}(x)\bigg]\hat{\psi}_{\mathrm{F}}(x)-mc\hat{\psi}_{\mathrm{F}}(x)$ $=0.$ (45)
When the coherent laser EM field is a plane-wave pulse, the Dirac field $\hat{\psi}_{\mathrm{F}}(x)$ admits an analytic solution in terms of the Volkov solutions [21, 31, 4, 6, 40, 20]. In this representation, the effect of the laser background is absorbed into the Dirac field itself, enabling scattering calculations to be performed with the laser influence incorporated from the outset.
@@ 5.1 Heisenberg-Euler vacuum model
$Z[\mathcal{A}]=e^{iW[\mathcal{A}]/\hbar}$ $=\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle,$ (46)
where, $W[\mathcal{A}]$ is the effective action whose integrand is the HE Lagrangian density [14, 6, 31, 36, 17, 39, 5]. Within the present approach, this expression can be reformulated in terms of a coherent state. We reconsider Eq. (8) in the Schrödinger picture. Suppose that the system contains only a coherent state $|\alpha_{\mathrm{S}}(t)\rangle$ at all times. We then examine the corresponding $\alpha$-$\alpha$ transition amplitude:
@@
Here, $C[\mathcal{A}]$ denotes the overall c-number phase generated by the action of the background-dressed free Hamiltonian on the vacuum $|0_{\mathrm{F}}\rangle$, i.e., the part of the Hamiltonian that governs the vacuum time evolution in the presence of the classical background but does not generate particle excitations, analogous to the vacuum energy phase in the Gell-Mann–Low theorem. This phase does not contribute to the Heisenberg–Euler effective action and drops out of all physical observables. A conceptual advantage of the coherent state formulation is that a background field automatically satisfies the Lorenz gauge condition $\partial_{\mu}\mathcal{A}^{\mu}(x)=0$ through the Gupta–Bleuler condition applied to the underlying photon coherent state.
The goal of this subsection is not to alter the value of the HE Lagrangian, but rather to clarify its physical origin within strong-field QED when classical background fields are understood as limits of quantum-coherent EM states. From the viewpoint of coherent-field QED, the physical interpretation of the HE vacuum becomes more transparent. In the Furry picture the quantity $\langle 0_{\mathrm{F}}|\hat{V}(\infty,-\infty)|0_{\mathrm{F}}\rangle$ is described as a “vacuum–vacuum transition” in the presence of an externally prescribed background field, a somewhat formal notion in which the vacuum is said to evolve non-trivially despite containing no photons. In contrast, the coherent-field QED formulation shows that the relevant transition is, more precisely, an $\alpha$-$\alpha$ transition: a coherent photon state representing a laser pulse propagates through the vacuum while retaining its quantum state. Nothing “happens” to the vacuum itself; rather, the effective action arises from the overall phase acquired by a coherent photon state propagating through the vacuum. This interpretation resolves the long-standing conceptual ambiguity of vacuum–vacuum transitions by clarifying that the HE effective action reflects the behavior of such a coherent state, rather than any dynamical evolution of an empty vacuum.
@@ 5.2 Beyond Furry picture
The present method also allows us to describe depletion effects of a coherent laser EM field. In the Furry picture, quantum transitions are treated under a fixed coherent state $\alpha$. Depletion effects correspond to quantum processes in which a coherent state $\alpha$ evolves into another coherent state $\alpha^{\prime}$. In the Schrödinger picture, the corresponding transition amplitude is
@@
An equivalent form can be found in the work of Ref. [6, 19]. The automatic inclusion of the Lorenz gauge condition is the advantage of the present method. In the present formulation, backreaction is not introduced as an independent mechanism, but appears automatically as the response of matter fields to transitions between coherent states (depletion). We note that depletion modeled via coherent-state deformations has been developed in [19]. Our treatment complements this by identifying the fixed-background approximation and its extensions as boundary-condition limits within a single first-principles framework.
@@ 6 Connection to BRST quantization
This section clarifies the relation between our coherent-state formulation and the BRST quantization of gauge theories [1, 38], ensuring the consistency of the physical-state conditions. The present formulation has been developed using the Gupta–Bleuler condition [15, 2], which provides a convenient and transparent implementation of gauge constraints in covariant QED. It is worth emphasizing, however, that this choice does not restrict the generality of the construction. In fact, the Gupta–Bleuler condition may be viewed as a special realization of the BRST formalism in the context of abelian gauge theory, where ghost degrees of freedom decouple from the physical sector (while the operator $\partial_{\mu}\hat{A}_{(+)\mathrm{H}}^{\mu}$ itself is not a Noether charge, the corresponding BRST charge provides the conserved generator that systematizes this condition into a cohomological definition of physical states).
@@
with equivalence classes taken modulo BRST-exact states. In QED, owing to the decoupling of ghost degrees of freedom, this condition reduces effectively to the Gupta–Bleuler constraint on the positive-frequency part of the gauge field [22, 26]. The coherent states considered in this work may therefore be equivalently characterized as BRST-physical states, as long as the displacement operator preserves the BRST cohomology.
@@
7 Coherent-state boundary conditions in path-integral representation
In contrast to conventional treatments of QED in prescribed background fields—where a classical field $\mathcal{A}$ is inserted directly into the action—our formulation shows that the same background-field functional integral arises as a boundary-condition limit of the full quantum theory, obtained by imposing coherent-state asymptotic conditions on the gauge field. This identifies the standard background-field path integral as a controlled limit of coherent-field QED rather than as an independent modification of the theory. The following derivation is standard except for the implementation of coherent-state boundary conditions, which encode the classical background field as a limiting case.
We reformulate the coherent-field QED amplitudes in the path-integral representation, showing explicitly how the standard generating functional of strong-field QED emerges from coherent-state boundary conditions. Throughout this section, $\mathcal{A}$ is treated as a classical field whose status depends on boundary conditions: it is fixed in the Furry picture limit, but becomes a dynamical integration variable when depletion is included. We rewrite the scattering amplitude
@@ 7 Coherent-state boundary conditions in path-integral representation
introduced in Eq. (48) as an overlap of asymptotic in–out states in the Heisenberg picture and translate it into a path-integral representation. Here, $\alpha_{(\mathrm{in})}$ and $\alpha_{(\mathrm{out})}^{\prime}$ denote coherent states of the EM field associated with the laser modes at $t\to$-$\infty$ and $t\to+\infty$, respectively. The states $\Psi_{(\mathrm{in})}$ and $\Psi_{(\mathrm{out})}^{\prime}$ represent general multiparticle states, including electrons, positrons and photons other than the laser modes. All states are understood as asymptotic in/out states in the Heisenberg picture. From the path-integral viewpoint, fixing coherent-state boundary conditions for the EM field yields the conventional generating functional of QED in a prescribed background field as a controlled limit, while allowing $\alpha\rightarrow\alpha^{\prime}$ transitions promotes the background to a dynamical variable integrated over histories consistent with the asymptotic coherent states. This makes explicit how the standard generating functional of strong-field QED and the Heisenberg–Euler effective action arise within our framework, without double counting of photon degrees of freedom [14, 36, 7, 6].
The starting point is the observation made in the depletion analysis [6, 19]: the amplitude $S_{fi}$ corresponds to a fully quantum transition between two distinct coherent photon states, accompanied by arbitrary changes in the remaining quantum degrees of freedom. This amplitude is therefore more general than the generating functional in a fixed background field, $Z[\mathcal{A}]$, employed in conventional strong-field QED, corresponding to the fixed-background limit of coherent-field QED which implicitly assumes a fixed background field and hence $\alpha^{\prime}=\alpha$.
To obtain a path-integral representation, recall Eq. (7), we separate the total gauge field into two components,
@@
$\mathcal{A}_{(-\infty)}^{\mu}(x)$ $=\langle\alpha_{(\mathrm{in})\mathrm{H}}|\hat{A}_{(\mathrm{in})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{in})\mathrm{H}}\rangle\quad\mathrm{for}\quad t\rightarrow-\infty,$ (52)
$\mathcal{A}_{(+\infty)}^{\mu}(x)$ $=\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}|\hat{A}_{(\mathrm{out})\mathrm{H}}^{\mu}(x)|\alpha_{(\mathrm{out})\mathrm{H}}^{\prime}\rangle\quad\mathrm{for}\quad t\rightarrow+\infty,$ (53)
for the asymptotic quantized EM field $\hat{A}_{(\mathrm{in/out})\mathrm{H}}^{\mu}(x)$. In contrast, the quantized EM field other than the laser field and the Dirac field are treated with standard in–out boundary conditions.
@@
$\hskip 0.0pt\langle\alpha_{(\mathrm{out})\mathrm{H}}^{\prime},\Psi_{(\mathrm{out})\mathrm{H}}^{\prime}|\alpha_{(\mathrm{in})\mathrm{H}},\Psi_{(\mathrm{in})\mathrm{H}}\rangle$ $=\int_{\mathcal{A}_{(-\infty)}}^{\mathcal{A}_{(+\infty)}}\mathcal{D}\mathcal{A}\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}$
@@
where the functional integration over $\mathcal{A}$ is restricted to field configurations consistent with the coherent state boundary conditions specified above. The Lagrangian density $\mathcal{L}_{\mathrm{QED}}$ is related in the standard way to the Hamiltonian density $\mathcal{H}_{\mathrm{QED}}$ employed in the previous operator-based formulation. This expression provides a systematic operator-based path integral formulation of laser depletion and backreaction, derived directly from operator-level matrix elements, without double counting of photon degrees of freedom.
The conventional generating functional of strong-field QED [6] is recovered as a special case. If the laser field is treated as fixed and undepleted, so that $\alpha^{\prime}=\alpha$ and the functional integration over $\mathcal{A}$ is suppressed, the above expression reduces to
$\hskip 0.0ptZ[\mathcal{A}]$ $=\int\mathcal{D}A\int\mathcal{D}\psi\int\mathcal{D}\bar{\psi}\exp\bigg[\frac{i}{\hbar c}\int d^{4}x\mathcal{L}_{\mathrm{QED}}(\psi(x),\bar{\psi}(x),A(x)+\mathcal{A}(x))\bigg]$ (55)
which is precisely the generating functional used in the Furry picture formulation and in the derivation of the Heisenberg–Euler effective action [14, 36]. From this viewpoint, $Z[\mathcal{A}]$ corresponds to a semiclassical limit of the more general coherent state transition amplitude, in which the laser field is frozen to its expectation value and quantum depletion effects are neglected.
We emphasize that the path-integral representation introduced here should be understood as a rewriting of operator-level matrix elements between physical (Gupta–Bleuler [15, 2]/BRST [1, 38, 22, 26]) states. Gauge constraints are therefore enforced prior to taking the functional integral, and the resulting expression does not represent an independent quantization procedure.